suppose potential energy between electron n proton at distance "r" is given by... (-ke^2)/(3.r^3)...use bohr's theory to obtain energy levels of such hypothetical atom

Grade:Upto college level

1 Answers

Chetan Mandayam Nayakar
12 years ago

V=potential energy,force of attraction=F=-dV/dr...(1), mvr=nh/(2pi)...(2),mv2/r=F...(3),E=K+U=(mv^2)/2 -c/r, where c is a universal constnt for electron-proton interaction, Solve equations (1),(2) and (3) to obtain the energy levels.